OpenStudy (anonymous):

What is the arc length if Θ = 2 pi over 7 and the radius is 3 cm? Please explain how to do it

OpenStudy (badhi):

the definition of the angle in radians is "the angle subtended at the centre of a circle by an arc equal in length to the radius" so if the arc length is D, \(D = r\theta\)

OpenStudy (campbell_st):

the arc length formula is \[l = r \times \theta\] l = arc length, substitute your information and claculate

OpenStudy (anonymous):

Thank you haha, that's all I need guys, but hey Camp ima give you a medal, can you give 1 to BAdhi for me? also what is the θ sign's name?

OpenStudy (campbell_st):

theta is the sign

OpenStudy (anonymous):

thanks :3